# Learning to Generate Instruction Tuning Datasets for Zero-Shot Task Adaptation

### v1

@@ Appendix C Generating Tasks with GPT-4
Here we use GPT-4 to generate tasks to adapt Mistral-7B-. We detail the process of generating synthetic instructing tuning datasets with GPT-4.
@@ C.2 Results
Appendix E Training Details
@@
E.1 Training Bonito
We train Mistral-7B on the conditional task generation with attributes (CTGA) dataset. From the training set, we uniformly sample 10,000 examples as the validation set to monitor the loss. The rest of the dataset is used for training Bonito. We train the model using Q-LoRA (Dettmers et al., 2023) by optimizing the cross entropy loss over the output tokens. The model is trained for 100,000 steps. The training takes about 4 days on four GPUs to complete. We include all the hyperparameters in Appendix E.5.
@@ E.1 Training Bonito
E.2 Instruction Tuned Models
Here we describe the procedure to train Mistral-7BP3 and Llama 2 7BP3. We use the processed T0 dataset from Muennighoff et al. (2022). Since the dataset is extremely large, we uniformly sample 1.6 million input-output examples and train the language model on them. Following Dettmers et al. (2023), we train the model for 10,000 steps with Q-LoRA and optimize the cross entropy loss over the output tokens. The training takes about 10 hours on four GPUs to complete. For the rest of the hyperparameters, see Appendix E.5.
@@ E.2 Instruction Tuned Models
E.3 Training Task-Specialized Models
@@
For fairness, we use the same hyperparameters to train task-specialized Bonito and the task-specialized Mistral-7B-Instruct-v0.2special and Mistral-7Bspecial models. Since the datasets have significantly fewer examples than CTGA, we train these models for at most 10,000 steps. If the training mixture has less than 160,000 examples, we train the Bonito model for 1 epoch. The training on four GPUs takes about 4 to 10 hours. For the rest of the hyperparameters, see Appendix E.5.
E.4 Software and Hardware Details
@@
E.5 Hyperparameters
Throughout our fine-tuning experiments, unless otherwise mentioned, we use the hyperparameters from Dettmers et al. (2023). Table 10 shows the hyperparameters in our experiments. We use gradient accumulation to achieve the effective batch size of 16. We also use gradient checkpointing which allows us to train large models like Llama 2 7B and Mistral-7B.
Appendix F Use of AI Assistants
@@
Appendix G Conditional Task Generation with Attributes: Datasets and Tasks
Table 11 shows the task distribution of the conditional task generation with attributes dataset. Table 12 lists all the datasets along with the task types in the dataset. The dataset includes 16 task types across 39 datasets. The task types are summarization, sentiment analysis, multiple-choice question answering, extractive question answering, topic classification, natural language inference, question generation, text generation, question answering without choices, paraphrase identification, sentence completion, yes-no question answering, word sense disambiguation, paraphrase generation, textual entailment, and coreference resolution. The difference between extractive question answering and question answering without choices is that in extractive question answering the target answer is present in the context whereas in question answering without choices, that always is not the case.
@@ Appendix G Conditional Task Generation with Attributes: Datasets and Tasks
Appendix H Prompts for Evaluation
H.1 PubmedQA
@@
Input
[…]
Answer Choices
Input
[…]
Answer Choices
Input
[…]
Answer Choices
Input
[…]
Answer Choices
Input
[…]
Answer Choices
H.2 Privacy Policy QA
@@
Input
Target
Answer Choices
Input
Target
Answer Choices
Input
Target
Answer Choices
Input
As a lawyer, can you answer the question given the context?Question: {{question}}Context:{{text}}Target
Answer Choices
Input
Target
Answer Choices
H.3 SQuADShifts
@@
H.3.1 NYT
Input
Target
Input
[…]
Input
[…]
Input
Target
Input
[…]
H.3.2 Amazon
Input
[…]
Input
[…]
Input
[…]
Input
Target
Input
[…]
H.3.3 Reddit
Input
[…]
Input
[…]
Input
[…]
Input
Target
Input
[…]
H.4 ContractNLI
@@
Input
Target
Answer Choices
Input
Target
Answer Choices
Input
Take the following as truth: {{premise}} Then the following statement: "{{hypothesis}}" is {{"true"}}, {{"false"}}, or {{"inconclusive"}}?Target
Answer Choices
Input
{{premise}} Based on that information, is the claim: "{{hypothesis}}" {{"true"}}, {{"false"}}, or {{"inconclusive"}}?Target
Answer Choices
Input
Target
Answer Choices
H.5 Vitamin C
@@
Input
Target
Answer Choices
Input
Target
Answer Choices
Input
Take the following as truth: {{evidence}} Then the following statement: "{{claim}}" is {{"true"}}, {{"false"}}, or {{"inconclusive"}}?Target
Answer Choices
Input
{{evidence}}Based on that information, is the claim: "{{claim}}" {{"true"}}, {{"false"}}, or {{"inconclusive"}}?Target
Answer Choices
Input
Target
Answer Choices
Appendix I Qualitatitve Examples
@@
Table 14: Example generations from Bonito for PubMedQA, SQuADShifts Amazon, and ContractNLI. -

### v2

@@ Appendix C Generating Tasks with GPT-4
Here we use GPT-4 to generate tasks to adapt Mistral-7BP3.
@@ C.2 Results
Appendix E Bonito with Smaller Models
We report an additional comparison with Bonito trained on Pythia (2.8B) (Biderman et al., 2023). We follow the same experimental setup used in Section 5.1.
| Yes-No QA | Extractive QA | NLI | ||||||||||
| Pythia (2.8B) | 23.7 0.0 | 42.2 1.4 | 11.9 0.9 | 8.9 0.5 | 8.0 0.6 | 20.8 3.5 | 25.4 1.5 | 20.1 | ||||
| Pythia (2.8B) + Bonito | 25.9 2.2 | 51.6 0.9 | 59.8 4.2 | 52.2 3.5 | 51.7 4.3 | 48.4 2.5 | 63.3 0.9 | 50.4 | ||||
Results
Table 11 shows that Bonito improves Pythia (2.8B) by an average of 30.3 F1 points across all the datasets. We observe that Pythia (2.8B) with Bonito performs better than Mistral with TAPT and Llama 2 with TAPT despite being twice as small (See Table 2). These results show that Bonito can be used to create small but powerful specialized language models.
Appendix F Training Details
@@
F.1 Training Bonito
We train Mistral-7B on the conditional task generation with attributes (CTGA) dataset. From the training set, we uniformly sample 10,000 examples as the validation set to monitor the loss. The rest of the dataset is used for training Bonito. We train the model using Q-LoRA (Dettmers et al., 2023) by optimizing the cross entropy loss over the output tokens. The model is trained for 100,000 steps. The training takes about 4 days on four GPUs to complete. We include all the hyperparameters in Appendix F.5.
@@ F.1 Training Bonito
F.2 Instruction Tuned Models
Here we describe the procedure to train Mistral-7BP3 and Llama 2 7BP3. We use the processed T0 dataset from Muennighoff et al. (2022). Since the dataset is large, we uniformly sample 1.6 million input-output examples and train the language model on them. Following Dettmers et al. (2023), we train the model for 10,000 steps with Q-LoRA and optimize the cross entropy loss over the output tokens. The training takes about 10 hours on four GPUs to complete. For the rest of the hyperparameters, see Appendix F.5.
@@ F.2 Instruction Tuned Models
F.3 Training Task-Specialized Models
@@
For fairness, we use the same hyperparameters to train task-specialized Bonito and the task-specialized Mistral-7B-Instruct-v0.2special and Mistral-7Bspecial models. Since the datasets have significantly fewer examples than CTGA, we train these models for at most 10,000 steps. If the training mixture has less than 160,000 examples, we train the Bonito model for 1 epoch. The training on four GPUs takes about 4 to 10 hours. For the rest of the hyperparameters, see Appendix F.5.
F.4 Software and Hardware Details
@@
F.5 Hyperparameters
Throughout our fine-tuning experiments, unless otherwise mentioned, we use the hyperparameters from Dettmers et al. (2023). Table 12 shows the hyperparameters in our experiments. We use gradient accumulation to achieve the effective batch size of 16. We also use gradient checkpointing to train large models like Llama 2 7B and Mistral-7B.
Appendix G Use of AI Assistants
@@
Appendix H Conditional Task Generation with Attributes: Datasets and Tasks
Table 13 shows the task distribution of the conditional task generation with attributes dataset. Table 14 lists all the datasets along with the task types in the dataset. The dataset includes 16 task types across 38 datasets. The task types are summarization, sentiment analysis, multiple-choice question answering, extractive question answering, topic classification, natural language inference, question generation, text generation, question answering without choices, paraphrase identification, sentence completion, yes-no question answering, word sense disambiguation, paraphrase generation, textual entailment, and coreference resolution. The difference between extractive question answering and question answering without choices is that in extractive question answering the target answer is present in the context whereas in question answering without choices, that always is not the case.
@@ Appendix H Conditional Task Generation with Attributes: Datasets and Tasks
Appendix I Prompts for Evaluation
I.1 PubmedQA
@@
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Given a passage: context.contexts | join(" ")
Answer the question: question
Summarize the above answer as YES, NO, or MAYBE?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django final_decision
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django yes ||| no ||| maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’m a doctor and I want to answer the question "question" using The following passage:
context.contexts | join(" ")
Summarize the above answer as YES, NO, or MAYBE?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django final_decision
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django yes ||| no ||| maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django What is the answer to the question "question" based on The following passage:
context.contexts | join(" ")
Summarize the above answer as YES, NO, or MAYBE?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django final_decision
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django yes ||| no ||| maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Please answer the question "question" using The following passage:
context.contexts | join(" ")
Summarize the above answer as YES, NO, or MAYBE?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django final_decision
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django yes ||| no ||| maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Given the following passage, answer the question: "question"
Passage: context.contexts | join(" ")
Summarize the above answer as YES, NO, or MAYBE?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django final_decision
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django yes ||| no ||| maybe
I.2 Privacy Policy QA
@@
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Given the context, is this related to the question? Context: text Question: question
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answer
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django Relevant|||Irrelevant
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Is this question "question" related to this context "text"?
Target {minted} [breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django Yes|||No
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Can this "text" help answer this question "question"?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django Yes|||No
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django As a lawyer, can you answer the question given the context? Question: question Context:text
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django Yes|||No
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Question:question Context:text Is the question related to the context?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django Yes|||No
I.3 SQuADShifts
@@
I.3.1 NYT
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django After reading the following paragraph, please answer this question: question
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’m working on the final exam for my class and am trying to figure out the answer to the question "question" I found the following info on New York Times and I think it has the answer. Can you tell me the answer?
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’ve always wondered: question
I searched New York Times and this is what I found. What’s the answer?
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django context
With the help of the passage, please answer the following question: question
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers["text"]|choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django ["Question", "Problem"] | choice range(1, 12) | choice: question
Hint: context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers["text"] | most_frequent | choice
I.3.2 Amazon
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django After reading the following paragraph, please answer this question: question
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’m working on the final exam for my class and am trying to figure out the answer to the question "question" I found the following info on Amazon and I think it has the answer. Can you tell me the answer?
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’ve always wondered: question
I searched Amazon and this is what I found. What’s the answer?
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django context
With the help of the passage, please answer the following question: question
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers["text"]|choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django ["Question", "Problem"] | choice range(1, 12) | choice: question
Hint: context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers["text"] | most_frequent | choice
I.3.3 Reddit
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django After reading the following paragraph, please answer this question: question
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’m working on the final exam for my class and am trying to figure out the answer to the question "question" I found the following info on Reddit and I think it has the answer. Can you tell me the answer?
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django I’ve always wondered: question
I searched Reddit and this is what I found. What’s the answer?
context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers[’text’] | most_frequent | choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django context
With the help of the passage, please answer the following question: question
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers["text"]|choice
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django ["Question", "Problem"] | choice range(1, 12) | choice: question
Hint: context
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answers["text"] | most_frequent | choice
I.4 ContractNLI
@@
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Suppose premise Can we infer that "hypothesis"? yes, no or maybe?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answer_choices[label]
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django No ||| Yes ||| Maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django premise
Question: Does this imply that "hypothesis"? yes, no or maybe?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answer_choices[label]
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django No ||| Yes ||| Maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Take the following as truth: premise Then the following statement: "hypothesis" is "true", "false", or "inconclusive"?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answer_choices[label]
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django False ||| True ||| Inconclusive
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django premise Based on that information, is the claim: "hypothesis" "true", "false", or "inconclusive"?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answer_choices[label]
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django False ||| True ||| Inconclusive
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django premise Based on the previous passage, is it true that "hypothesis"? Yes, no, or maybe?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django answer_choices[label]
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django No ||| Yes ||| Maybe
I.5 Vitamin C
@@
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Suppose evidence Can we infer that "claim"? yes, no or maybe?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django No ||| Yes ||| Maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django evidence
Question: Does this imply that "claim"? yes, no or maybe?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django No ||| Yes ||| Maybe
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django Take the following as truth: evidence Then the following statement: "claim" is "true", "false", or "inconclusive"?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django False ||| True ||| Inconclusive
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django evidence Based on that information, is the claim: "claim" "true", "false", or "inconclusive"?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django False ||| True ||| Inconclusive
Input {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django evidence Based on the previous passage, is it true that "claim"? Yes, no, or maybe?
Target {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=,bgcolor=bg]django
Answer Choices {minted}[breaklines, tabsize=2,breaksymbolleft=, fontsize=, bgcolor=bg]django No ||| Yes ||| Maybe
Appendix J Qualitatitve Examples
@@
Table 16: Example generations from Bonito for PubMedQA, SQuADShifts Amazon, and ContractNLI.